OpenStudy (anonymous):

domain of square root of 5x-20

OpenStudy (yttrium):

The domain of this equation can be expressed as \[\sqrt{5x-20} \ge0\] Therefore, @sbarthole ?

OpenStudy (anonymous):

It's a tautology to say \[ \sqrt{f(x)}\geq 0 \]When \(f(x)\in\mathbb R\).

OpenStudy (anonymous):

It's a positive square root. What you wanna check is that \[ 5x-20\geq 0 \]

OpenStudy (yttrium):

Thanks for emphasizing @wio.